OpenStudy (anonymous):

What is the y-intercept and the x-intercept (ordered pairs) for the following equation: -10x + y = 6

OpenStudy (anonymous):

y int=(0,6) x int= ( -3/5,0)

OpenStudy (anonymous):

First convert it to slope-intercept form y=mx+b (y=10x+6) To find your y intercept, you set the x in the equation equal to zero. So: y=10(0)+6 y=6 X-intercept is the same way, just set your y to zero: 0=10x+6 -6=10x -6/10=x (Simplify fractions) x=-3/5